# RepoExec: Evaluate Code Generation with a Repository-Level Executable Benchmark

### v2

@@ Abstract
The ability of CodeLLMs to generate executable and functionally correct code at the repository-level scale remains largely unexplored. We introduce RepoExec, a novel benchmark for evaluating code generation at the repository-level scale. RepoExec focuses on three main aspects: executability, functional correctness through automated test case generation with high coverage rate, and carefully crafted cross-file contexts to accurately generate code. Our work explores a controlled scenario where developers specify necessary code dependencies, challenging the model to integrate these accurately. Experiments show that while pretrained LLMs outperform instruction-tuned models in correctness, the latter excel in utilizing provided dependencies and demonstrating debugging capabilities. We also introduce a new instruction-tuned dataset that focuses on code dependencies and demonstrate that CodeLLMs fine-tuned on our dataset have a better capability to leverage these dependencies effectively. RepoExec aims to provide a comprehensive evaluation of code functionality and alignment with developer intent, paving the way for more reliable and applicable CodeLLMs in real-world scenarios. The dataset and source code can be found at https://github.com/FSoft-AI4Code/RepoExec.
@@
1 Introduction
Code Large Language Models (CodeLLMs) have emerged as a prominent solution to assist with coding tasks recently Wang et al. (2021, 2023); Feng et al. (2020); Allal et al. (2023); Li et al. (2023); Lozhkov et al. (2024); Guo et al. (2024); Pinnaparaju et al. (2024); Zheng et al. (2024); Roziere et al. (2023); Nijkamp et al. (2022); Luo et al. (2023); Xu et al. (2022); Bui et al. (2023). While CodeLLMs are adept at capturing the immediate context of the code being written, they often do not fully exploit the broader context available within the entire code repository. This limitation stems from the token length restrictions of general LLMs, which also apply to CodeLLMs. Consequently, suggestions generated by these models may not be optimally aligned with the project’s architecture or intended functionality, as they tend to overlook the rich information embedded in related files, class hierarchies, dependencies, and more. Therefore, evaluating the ability of CodeLLMs on code generation at the repository-level scale is crucial to accurately measure their performance in real-world scenarios. There are existing repository-level code generation benchmarks , such as RepoBench Liu et al. (2023b), RepoCoder Zhang et al. (2023b), CrossCodeEval Ding et al. (2023), CoCoMIC Ding et al. (2024), and DevEval Li et al. (2024), Although these benchmarks provide a comprehensive perspective on evaluating code generation at the repository-level scale, they lack 3 major features that prevent them from accurately reflecting real-world coding scenarios. First, code generation in the real world requires the code to be executable, and none of the existing benchmarks focus on this aspect. Second, these benchmarks rely on simple methods to retrieve contexts, such as including all imports and function calls (RepoBench) or simply retrieving similar files as contexts (RepoCoder), which might introduce noise into the contexts and impact the performance of code generation systems on these benchmarks. Third, generating real world code at repository-level scale requires the CodeLLMs to be able to capture contexts (both cross-file and in-file effectively), which may pose several questions such as if the dependencies provided by human are really the dependencies that the LLM need, or how well the CodeLLMs process the dependencies given that lots of depdencies feed into the prompt, do they use all of them or skip some of them?
@@ 1 Introduction
To effectively address these shortcomings, we introduce RepoExec, a novel benchmark for evaluating code generation at the repository level that emphasizes executability and correctness. This benchmark is specifically designed to overcome the limitations of existing benchmarks by focusing on real-world applicability and a comprehensive evaluation of code functionality with 3 main focus:
  • Enhanced Executability: RepoExec ensures that generated code is not only syntactically correct but also fully executable in a real-world environment by establishing an automated system that verifies installation and runtime requirements.
@@
  • Dynamic Test Case Generation: RepoExec incorporates a sophisticated mechanism for dynamically generating high-coverage test cases tailored to the functionality of the newly generated code, ensuring that the code performs its intended tasks accurately.
@@
In addition, we introduce the Dependency Invocation Rate (DIR), a novel metric that measures the proportion of code dependencies successfully integrated into the generated code. Higher DIR values indicate better model comprehension of dependencies, while lower values suggest challenges in dependency utilization.
Our experiments with RepoExec reveal key insights into repository-level code generation by CodeLLMs. Models achieve optimal performance with full context dependencies, though smaller contexts outperform medium ones due to formatting issues in the BasePrompt. While pretrained LLMs like Codellama-34b-Python lead in pass@1 rates, instruction-tuned LLMs is better at managing dependencies, despite sometimes generating overly complex code. In contrast, pretrained LLMs produce accurate code but struggle with dependency utilization. Multi-round debugging tests, especially with GPT-3.5 and WizardCoder, significantly improve both pass@1 scores and DIR after three rounds. Fine-tuning with our dependency-enhanced dataset not only boosts pass@1 and DIR metrics but also minimizes computational costs, highlighting the importance of executable code testing and the advantages of instruction-tuning and multi-round debugging in enhancing CodeLLMs’ performance and dependency management. In summary, our contributions are as follows:
@@
    We introduce RepoExec, a novel benchmark designed to evaluate code generation at the repository level with a focus on executability and correctness. This benchmark addresses the gaps in existing systems by emphasizing real-world applicability and providing a comprehensive assessment of code functionality.
@@
    RepoExec ensures that generated code is not only syntactically correct but also fully executable in real-world environments. This is achieved through an automated system that verifies installation and runtime requirements, ensuring the code can function in actual software settings.
@@
    Our benchmark includes a sophisticated mechanism for dynamically generating high-coverage test cases that are tailored to the functionality of the newly generated code. This ensures the code not only runs but also accurately performs its intended tasks.
@@
    We implement a pipeline within RepoExec to assess how effectively CodeLLMs utilize code dependencies. This includes an novel instruction-tuning dataset that enriches CodeLLMs’ capability to leverage these dependencies for improved code generation.
@@
    We introduce the Dependency Invocation Rate (DIR), a novel metric that measures the proportion of provided dependencies successfully incorporated into the generated code. This metric helps gauge the models’ understanding and utilization of dependencies.
@@
    Our experiments provide crucial insights into the performance of CodeLLMs in repository-level code generation. We demonstrate that while foundation models achieve high initial accuracy, instruction-tuned models excel in dependency management and addressing complex scenarios. Additionally, our multi-round debugging tests show notable improvements in model performance, particularly with enhancements in dependency management.
@@ 2 Related works
Coding-related tasks have been crucial for assessing the performance of Large Language Models (LLMs), with code generation emerging as a primary focus (Chen et al., 2021; Li et al., 2023; Jiang et al., 2024; Touvron et al., 2023; Roziere et al., 2023; Xu et al., 2022; Allal et al., 2023; Nijkamp et al., 2022). Early benchmarks have been introduced to address this issue Yin et al. (2018); Iyer et al. (2018); Manh et al. (2023); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); however, they often had limited scope or employed weak evaluation approaches. Some benchmarks Yin et al. (2018); Iyer et al. (2018); Manh et al. (2023) exhibit domain diversity akin to real-world applications; however, their evaluation methodologies are constrained to match-based metrics, thereby decreasing the reliability of these benchmarks Chen et al. (2021). Meanwhile, benchmarks with reliable evaluation approaches such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021) and APPS Hendrycks et al. (2021) often entail limitations to specific domains like competitive programming. Besides, all the benchmarks mentioned primarily focus on standalone function generation, lacking consideration for cross-contextual and dependency invocation scenarios.
@@ 3.1 Data Source
Developing an executable benchmark within repository contexts is challenging due to complex setup requirements and frequent lack of clear installation guidelines in repositories. Previous studies (Ding et al., 2024; Shrivastava et al., 2023; Ding et al., 2023; Liao et al., 2023) have often relied on match-based metrics for evaluation, which may not fully capture code functionality. In addition, test cases are essential for assessing code functionality. However, extracting test cases from repositories (Zhang et al., 2023a; Li et al., 2024; Zhang et al., 2024) often relies on available functions and heuristic rules, limiting adaptability and excluding data without existing tests. For instance, Li et al. 2024 found that over 99% of functions were discarded due to the absence of suitable tests. We propose a dataset collection pipeline to ensure that repositories can build executable environments and that test cases are automatically generated.
@@ 3.2 Functions and Dependencies Extraction
We employ static analysis to identify dependencies, excluding identifiers that are function parameters or typing objects. Each dependency name is then mapped to its definition within the repository using a repository graph and static analysis tools.
@@ 3.3 Test case generation
In our test generation process, we conduct two phases corresponding to Correctness Control and Coverage Enhancement. After the test generation process, we exclude samples with a coverage percentage lower than 40%, as they are of insufficient quality to accurately assess the correctness of the generated code.
@@ 3.3.2 Coverage Enhancement
Weak unit tests may allow incorrect implementations to pass Liu et al. (2023a). To address this, we propose a strategy for enhancing test coverage using LLMs. Given the complexity of this task, requiring a strong understanding of code, we utilize GPT-3.5 to improve the quality of test cases. We provide GPT-3.5 with three prompts (see Appendix C) to handle challenging scenarios, including edge and corner cases. The initial generated tests serve as few-shot examples. We then extract the newly generated tests and ensure their correctness using our methodology from Section 3.3.1. Table 1 shows a coverage improvement of about 4% after enhancement, reaching 96.25%. The performance gap has also increased to over 5% (Appendix B), indicating greater robustness in the evaluation.
@@ 4.2 Dataset Stastistic
Dataset #Samples Repo context Test case Coverage (%) CoNaLA Yin et al. (2018) 500 - CONCODE Iyer et al. (2018) 2,000 - HumanEval Chen et al. (2021) 164 ✓, H 99.43 MBPP Austin et al. (2021) 974 ✓, H 98.48 RepoCoder Zhang et al. (2023a) 373 ✓, P - CrossCodeEval Ding et al. (2023) 2,665 - CoCoMIC Ding et al. (2024) 6,888 - DevEval Li et al. (2024) 1,874 ✓, P - RepoExec 355 ✓, A 92.46    + coverage-enhancement 96.25
Table 1: The comparison between popular code generation benchmarks and RepoExec. For test case, we denote H: Human annotated, P: Pre-existing, A: Automated.
Table 1 compares the details of RepoExec with those of existing code generation benchmarks. Benchmarks that exclude execution-based evaluation Yin et al. (2018); Iyer et al. (2018); Ding et al. (2023, 2024) may gather substantial amounts of data; however, they are inadequate for assessing the quality of the generated code. For HumanEval and MBPP, the majority of problems involve standalone functions, which do not reflect real-world scenarios. Besides, benchmarks that rely on human-annotated and pre-existing test cases Chen et al. (2021); Austin et al. (2021); Zhang et al. (2023b); Li et al. (2024) are challenging to scale and control the test coverage. Finally, repository-context benchmarks (RepoCoder Zhang et al. (2023b), CrossCodeEval Ding et al. (2023), CoCoMIC Ding et al. (2024), and DevEval Li et al. (2024)) primarily focus on evaluating retrieval modules. In contrast, our approach emphasizes the generation module’s ability to understand and utilize human-provided dependency contexts.
@@
#No problem #No testcase Avg tokens
Cross-file Total Avg Coverage (%) Avg Prompt Solution
Full context 22.8% 355 96.25 98.38 362.92 78.46
Medium context - - - - 253.05 -
Short context - - - - 179.66 -
Table 2: Dataset attributes. Cross-file refers to the percentage of problems that involve cross-file dependencies. Tree-sitter is used for tokenization.

### v3

# On the Impacts of Contexts on Repository-Level Code Generation

@@ Abstract
CodeLLMs have gained widespread adoption for code generation tasks, yet their capacity to handle repository-level code generation with complex contextual dependencies remains underexplored. Our work underscores the critical importance of leveraging repository-level contexts to generate executable and functionally correct code. We present RepoExec, a novel benchmark designed to evaluate repository-level code generation, with a focus on three key aspects: executability, functional correctness through comprehensive test case generation, and accurate utilization of cross-file contexts. Our study examines a controlled scenario where developers specify essential code dependencies (contexts), challenging models to integrate them effectively. Additionally, we introduce an instruction-tuned dataset that enhances CodeLLMs’ ability to leverage dependencies, along with a new metric, Dependency Invocation Rate (DIR), to quantify context utilization. Experimental results reveal that while pretrained LLMs demonstrate superior performance in terms of correctness, instruction-tuned models excel in context utilization and debugging capabilities. RepoExec offers a comprehensive evaluation framework for assessing code functionality and alignment with developer intent, thereby advancing the development of more reliable CodeLLMs for real-world applications. The dataset and source code are available at https://github.com/FSoft-AI4Code/RepoExec.
1 Introduction
@@
Code Large Language Models (CodeLLMs) have emerged as powerful tools for assisting with coding tasks Wang et al. (2021, 2023); Feng et al. (2020); Allal et al. (2023); Li et al. (2023); Lozhkov et al. (2024); Guo et al. (2024); Pinnaparaju et al. (2024); Zheng et al. (2024); Roziere et al. (2023); Nijkamp et al. (2022); Luo et al. (2023); Xu et al. (2022); Bui et al. (2023). While these models excel at generating code from natural language requirements or completing individual lines of code, their application in real-world, professional software development scenarios presents more complex challenges. A critical aspect of this complexity lies in leveraging relevant contexts across an entire software repository, which raises two pivotal questions. First, to what extent are the retrieved dependencies accurate and relevant, rather than potential noise in the input? Second, how effectively do LLMs process and incorporate the provided dependencies into their generated code? These inquiries are central to understanding the capabilities and limitations of CodeLLMs in repository-level code generation, where completing a single line of code might require making API calls to functions within the same file (in-file context) or across different files (cross-file context).
Existing repository-level code generation benchmarks, such as RepoBench Liu et al. (2023b), RepoCoder Zhang et al. (2023a), CrossCodeEval Ding et al. (2023), CoCoMIC Ding et al. (2024), and DevEval Li et al. (2024), have made significant strides in assessing code generation at a repository-level scale. However, these benchmarks face several limitations that hinder their ability to comprehensively evaluate real-world coding scenarios:
  1. 1.
    Lack of an executable environment: This leads to reliance on match-based metrics, which are not robust for assessing the functional correctness of generated code.
  2. 2.
    Inadequate control over test quality: Although RepoCoder and DevEval ensure executability, they depend on pre-existing test cases extracted from repositories, limiting the robustness of evaluation and hindering the scalability of the data pipeline Liu et al. (2023a).
  3. 3.
    Overreliance on functional correctness metrics like pass@k: These metrics, inherited from standalone code generation studies Chen et al. (2021); Austin et al. (2021), are insufficient in the context of repository-level code generation.
Moreover, real-world development practices often involve breaking down large functions into smaller modules (functions, classes, or variables) to enhance problem-solving efficiency. Consequently, generated code must effectively call these predefined dependencies to align with developer intent. However, LLMs may produce code that passes test cases but relies on inefficient workarounds or unnecessarily reimplements predefined functions. This can lead to suboptimal implementations, code duplication, and potential risks such as technical debt and code smell issues (Maldonado and Shihab, 2015; Sierra et al., 2019; Rasool and Arshad, 2015; Santos et al., 2018).
Given these challenges, it is evident that evaluating CodeLLMs based solely on functional correctness is insufficient. There is a pressing need for a more comprehensive evaluation strategy that assesses not only the correctness of generated code but also its alignment with developer intent, efficiency in utilizing provided dependencies, and adherence to best practices in software development. Crucially, our research reveals a strong correlation between functional correctness (as measured by pass@1) and effective dependency utilization, suggesting that models that better leverage contextual information tend to produce more accurate and functionally correct code.
To effectively address these shortcomings, we introduce a paradigm for evaluation that emphasizes both functional correctness and dependency utilization, while also exploring their relation. Specifically, we introduce the Dependency Invocation Rate (DIR), a novel metric that measures the proportion of code dependencies successfully integrated into the generated code. Higher DIR values indicate better model comprehension of dependencies, while lower values suggest challenges in dependency utilization. Importantly, our findings demonstrate a strong correlation between DIR and pass@1 rates, underscoring the critical role of effective context utilization in producing high-quality, functionally correct code.
@@
We also introduce RepoExec, a novel benchmark specifically designed to overcome the limitations of existing benchmarks, providing a robust and comprehensive evaluation for code generation at real-world applicability. In detail, RepoExec advances with two primary focuses:
@@
  • Executability and Dynamic Test Case Generation: RepoExec ensures an executable environment within the repository context, thereby providing a reliable evaluation of functional correctness. Besides, it also incorporates a practical mechanism for dynamically generating high-coverage test cases tailored to the functionality of the newly generated code, ensuring that the code performs its intended tasks accurately.
@@
Our experiments with RepoExec reveal key insights into repository-level code generation by CodeLLMs. Models achieve optimal performance with full context dependencies, though smaller contexts outperform medium ones due to formatting issues in the BasePrompt. While pretrained LLMs like Codellama-34b-Python lead in pass@1 rates, instruction-tuned LLMs are better at managing dependencies, despite sometimes generating overly complex code. In contrast, pretrained LLMs produce accurate code but struggle with dependency utilization. These findings further reinforce the observed correlation between pass@1 and DIR, highlighting the importance of balancing accuracy and context utilization in CodeLLM performance.
Furthermore, we explore two potential approaches for enhancing the quality of generated code: Multi-round debugging and Instruction tuning. Multi-round debugging with test execution, especially with GPT-3.5 and WizardCoder, significantly improves both pass@1 scores and DIR after three rounds. Fine-tuning with our dependency-enhanced dataset not only boosts pass@1 and DIR metrics but also minimizes computational costs, highlighting the importance of executable code testing and the advantages of instruction-tuning and multi-round debugging in enhancing CodeLLMs’ performance and dependency management.
In summary, our contributions are as follows:
@@
    We introduce a novel evaluation paradigm aimed at providing a robust and comprehensive assessment for repository-level code generation. Our approach not only evaluates the functional correctness of generated code but also considers additional quality factors such as maintainability and adherence to clean code principles through effective dependency utilization.
@@
    We present the Dependency Invocation Rate (DIR), a novel metric that measures the proportion of provided dependencies successfully incorporated into the generated code. This metric helps gauge the models’ understanding and utilization of dependencies and shows a strong correlation with functional correctness.
@@
    We introduce RepoExec, a novel benchmark that aligns with our evaluation paradigm, addressing the gaps in existing benchmarks and offering a comprehensive assessment of code generation quality.
@@
    We implement an effective pipeline within RepoExec from dependency extraction, dynamically generate high-coverage test cases and automatic evaluation with execution and code dependencies. Our pipeline offers practical usage and scalability for the community.
@@
    We release a tool named pydepcall [^1] https://github.com/FSoft-AI4Code/pydep/tree/main to extract dependencies for all functions within any repository, providing a practical use for advancing research in this domain.
@@
    Our experiments provide crucial insights into the performance of CodeLLMs in repository-level code generation. We demonstrate that while foundation models achieve high initial accuracy, instruction-tuned models excel in dependency management and addressing complex scenarios. Additionally, our multi-round debugging tests show notable improvements in model performance, particularly with enhancements in dependency management. Importantly, we establish a strong correlation between pass@1 and DIR, emphasizing that models which better utilize contexts tend to produce more functionally correct code.
@@ 2 Related works
Coding-related tasks have been crucial for assessing the performance of Large Language Models (LLMs), with code generation emerging as a primary focus (Chen et al., 2021; Li et al., 2023; Jiang et al., 2024; Touvron et al., 2023; Roziere et al., 2023; Xu et al., 2022; Allal et al., 2023; Nijkamp et al., 2022). Early benchmarks have been introduced to address this issue Yin et al. (2018); Iyer et al. (2018); Nguyen et al. (2023); Chen et al. (2021); Austin et al. (2021); Hendrycks et al. (2021); however, they often had limited scope or employed weak evaluation approaches. Some benchmarks Yin et al. (2018); Iyer et al. (2018); Nguyen et al. (2023) exhibit domain diversity akin to real-world applications; however, their evaluation methodologies are constrained to match-based metrics, thereby decreasing the reliability of these benchmarks Chen et al. (2021). Meanwhile, benchmarks with reliable evaluation approaches such as HumanEval Chen et al. (2021), MBPP Austin et al. (2021) and APPS Hendrycks et al. (2021) often entail limitations to specific domains like competitive programming. Recently, there have been efforts to extend the domains of generation tasks in various benchmarks. For instance, ExeDS Huang et al. (2022) focuses on data science code generation, while ODEX Wang et al. (2022) serves as an open-domain benchmark for code generation. Besides, all the benchmarks mentioned primarily focus on standalone function generation, lacking consideration for cross-contextual and dependency invocation scenarios.
@@
3 Evaluation Paradigm
In this section, we outline our approach to achieve a more robust and comprehensive evaluation of repository-level code generation. Our paradigm encompasses two key attributes: Functional correctness and Dependency utilization.
Functional correctness: This evaluation criterion ensures that the code accurately performs its intended tasks and requirements. Specifically, it typically involves using test cases to compare the execution output of the generated code against the expected output for a given input. This criterion has been widely employed for evaluating code generation in numerous studies (Chen et al., 2021; Austin et al., 2021; Zhuo et al., 2024; Li et al., 2024). We follow the well-known automatic metric Pass@k Chen et al. (2021) to measure the functional correctness of generation outputs.
Dependency utilization: However, functional correctness alone cannot fully assess code quality, as it may overlook issues like poor implementation, workaround solutions, or code redundancy. Assuming that human-written solutions are optimal, match-based metrics like BLEU and edit similarity can assess the alignment between the generated code and the high-quality reference solution at the token, word, or character level. However, not all words and tokens contribute to the quality of the code and alternative implementations can preserve the quality despite low similarity scores. For instance, local variable names or comments can vary widely without impacting code quality. Additionally, statements like while and for can be interchangeable while preserving the algorithm’s complexity. Thus, these match-based metrics are also not reliable for accurately measuring code quality. Meanwhile, tokens representing called dependencies—such as third-party packages, functions, variables, and classes within the repository—demonstrate effective use of quality code for efficient implementation. In contrast, ignoring these dependencies may suggest workaround implementations misaligned with human intent, leading to extensive verification and maintenance costs.
Therefore, to assess the models’ ability to utilize provided dependencies in accordance with human intent, we introduce the Dependency Invocation Rate (DIR). This metric represents the percentage of invoked dependencies out of the total number of dependencies provided. Let $D_{g}$ denote the set of identifiers in the generated output, and $D_{s}$ denote the set of provided dependencies extracted from the solution. The Dependency Invocation Rate (DIR) is calculated as follows:
$\text{DIR}=\dfrac{|D_{g}\cap D_{s}|}{|D_{s}|}$
In summary, we state that achieving a high-quality solution requires the generated code to excel in both functional correctness and dependency utilization. Otherwise, it indicates a lack of ability to generate correct solutions for the intended tasks or suggests poor implementation practices, including technical debt or code smell issues.
@@ 4.1 Data Source
Developing an executable benchmark within repository contexts is challenging due to complex setup requirements and frequent lack of clear installation guidelines in repositories. Previous studies (Ding et al., 2024; Shrivastava et al., 2023; Ding et al., 2023; Liao et al., 2023) have often relied on match-based metrics for evaluation, which may not fully capture code functionality. In addition, test cases are essential for assessing code functionality. However, extracting test cases from repositories (Zhang et al., 2023b; Li et al., 2024; Zhang et al., 2024) often relies on available functions and heuristic rules, limiting adaptability and excluding data without existing tests. For instance, Li et al. 2024 found that over 99% of functions were discarded due to the absence of suitable tests. We propose a dataset collection pipeline to ensure that repositories can build executable environments and that test cases are automatically generated.
@@ 4.2 Functions and Dependencies Extraction
We employ static analysis to identify dependencies, excluding identifiers that are function parameters or typing objects. Each dependency name is then mapped to its definition within the repository using a repository graph and static analysis tools. We release our tool pydepcall for community usage and provide a brief description in Appendix H.
@@ 4.3 Test case generation
In our test generation process, we conduct two phases corresponding to Correctness Control and Coverage Enhancement. After the test generation process, we exclude samples with a line coverage lower than 40%, as they are of insufficient quality to accurately assess the correctness of the generated code.
@@ 4.3.2 Coverage Enhancement
Weak unit tests may allow incorrect implementations to pass Liu et al. (2023a). To address this, we propose a strategy for enhancing test coverage using LLMs. Given the complexity of this task, requiring a strong understanding of code, we utilize GPT-3.5 to improve the quality of test cases. We provide GPT-3.5 with three prompts (see Appendix C) to handle challenging scenarios, including edge and corner cases. The initial generated tests serve as few-shot examples. We then extract the newly generated tests and ensure their correctness using our methodology from Section 4.3.1. Table 1 shows a line coverage improvement of about 4% after enhancement, reaching 96.25%. The performance gap has also increased to over 5% (Appendix B), indicating greater robustness in the evaluation.
@@ 5.2 Dataset Stastistic
Dataset #Samples Context TC LC (%) CoNaLA Yin et al. (2018) 500 - CONCODE Iyer et al. (2018) 2,000 - HumanEval Chen et al. (2021) 164 ✓, H 99.43 MBPP Austin et al. (2021) 974 ✓, H 98.48 RepoCoder Zhang et al. (2023b) 373 ✓, P - CrossCodeEval Ding et al. (2023) 2,665 - CoCoMIC Ding et al. (2024) 6,888 - DevEval Li et al. (2024) 1,874 ✓, P - RepoExec 355 ✓, A 92.46    + coverage-enhancement 96.25
Table 1: The comparison between popular code generation benchmarks and RepoExec. For test case, we denote H: Human annotated, P: Pre-existing, A: Automated. TC: Test Cases. LC: Line Coverage
Table 1 compares the details of RepoExec with those of existing code generation benchmarks. Benchmarks that exclude execution-based evaluation Yin et al. (2018); Iyer et al. (2018); Ding et al. (2023, 2024) may gather substantial amounts of data; however, they are inadequate for assessing the quality of the generated code. For HumanEval and MBPP, the majority of problems involve standalone functions, which do not reflect real-world scenarios. Besides, benchmarks that rely on human-annotated and pre-existing test cases Chen et al. (2021); Austin et al. (2021); Zhang et al. (2023a); Li et al. (2024) are challenging to scale and control the test coverage. Finally, repository-context benchmarks (RepoCoder Zhang et al. (2023a), CrossCodeEval Ding et al. (2023), CoCoMIC Ding et al. (2024), and DevEval Li et al. (2024)) primarily focus on evaluating retrieval modules. For example, RepoCoder introduces a retriever using a sparse bag-of-words model and provides the effectiveness of this retriever. Similarly, CrossCodeEval experiments and reports performance using various retrievers such as BM25, UniXCoder, and OpenAI ada. CoCoMic proposed CCFINDER to retrieve cross-file context. DevEval is the closest to our work; however, they compare the performance of the generation model based on different given types of context, which can also align with the different contexts provided by different retrievers. In contrast, our approach emphasizes the generation module’s ability to understand and utilize human-provided dependency contexts.
@@
#No problem #No testcase Avg tokens Cross-file Total Avg LC (%) Avg Prompt Solution Full 22.8% 355 96.25 99.45 362.92 78.46 Medium - - - - 253.05 - Small - - - - 179.66 -
Table 2: Dataset attributes with different levels of contexts (Full, Medium, Small). Cross-file refers to the percentage of problems that involve cross-file dependencies. Tree-sitter is used for tokenization. AVG LC: Average Line Coverage.